OpenStudy (anonymous):

Evalute the limit: lim 4-(12+x)^(1/2) x->4 (x-4)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 4} (4 - \sqrt{12+x})/(x-4)\]

OpenStudy (bahrom7893):

multiply both top and bottom by 4+sqrt(12+x)

OpenStudy (anonymous):

I did and I (16-16) / 4-4

OpenStudy (anonymous):

And I got*

OpenStudy (bahrom7893):

(16-(12+x))/[(x-4)(4+sqrt(12+x))] (16-12-x)/[(x-4)(4+sqrt(12+x))]

OpenStudy (bahrom7893):

-(x-4)/[(x-4)(4+sqrt(12+x))] -1/(4+sqrt(12+x)) Plug in x=4 -1/(4+sqrt(12+4)) -1(4+sqrt(16)) -1/(4+4) -1/8

OpenStudy (anonymous):

Thank you so mmuch :) I applied the limit, and forgot to rearrange 4-x to -(x-4)